# DEAR: Dataset for Evaluating the Aesthetics of Rendering

### v3

@@ Abstract
Figure 1: Limitations of existing IQA methods: while LIQE [Zhang2023] assigns nearly identical quality scores with standard deviations from the mean value 3.5% across evaluations, subjective evaluation reveals clear differences. Our work addresses this gap by introducing a new benchmark Dataset for Evaluating of Aesthetics of Rendering (DEAR).
@@ 1 Introduction
Image Quality Assessment refers to the task of automatically evaluating the quality of images in a way that aligns with human perception. Accurate IQA is critical for the development and optimization of computer vision and image processing algorithms, as it provides an objective signal that guides models toward producing results that are visually pleasing or perceptually faithful [zhai2020perceptual]. Whether in image compression, enhancement, restoration, or generation, IQA metrics are expected to help ensure that algorithmic outputs meet human expectations of quality, as well as to measure the performance of the algorithms themselves.
Traditional IQA research has focused primarily on assessing the degradation of perceived image quality introduced by the presence of distortions such as noise, blur, or compression artifacts. Over time, however, the field recognized that perceived image quality extends beyond simple degradations: viewers judge an image not only by the absence of artifacts, but also by how convincingly it conveys realism and visual appeal. This shift in perspective motivated the creation of modern subjective IQA datasets, such as UHD-IQA [Hosu2024], which capture human judgments on more holistic attributes, including composition, resolution, and color harmony, without requiring a pristine reference. However, even in subjective IQA datasets [Hosu2024], participants are typically asked to rate the overall quality of individual images, not to compare different stylistic renditions of the same scene. As a result, such datasets capture judgments about technical fidelity or perceptual realism rather than aesthetics of rendering, and are not suited for modeling user preferences between alternative rendering interpretations, which could be captured much more easily if datasets such as UHD-IQA contained pairwise evaluation results.
Evaluation of aesthetics of rendering introduces a qualitatively different challenge. Aesthetics of rendering is not characterized by degradations, but by aesthetic attributes such as color usage, texture, and artistic expression. Importantly, while image degradations are known to be of lesser quality than the original image, multiple renderings can be equally valid and pleasing, making rendering preference inherently subjective and context-dependent. Unlike quality degradations, which tend to have a broadly agreed-upon negative impact, rendering variation often reflects personal, cultural, or situational preferences. Finally, because there is no objectively correct answer for the most pleasant rendering, it is impossible to guarantee whether users select their responses attentively, making the annotation process vulnerable to careless behavior [ershov2024reliability].
@@
To address this gap, we introduce the Dataset for Evaluating the Aesthetics of Rendering (DEAR), a new benchmark specifically designed to allow learning subjective judgments of image rendering. DEAR is constructed by augmenting the MIT-Adobe FiveK dataset [fivek] with pairwise human preference scores obtained through crowdsourced evaluation platform. Each pair of image renderings was evaluated by at least $25$ evaluators, with a total of $13,648$ unique evaluators participating overall. These annotations enable the development of EAR models that go beyond degradation-focused assessments and instead capture the more nuanced, human-centered perception of rendering.
@@ 2.1 Benchmark IQA datasets
In recent years, numerous IQA datasets have been introduced. Benchmarks such as LIVE [sheikh2005live], VCL@FER [Zari2012], NITS-IQA [Ruikar2023], and CSIQ [Chandler2010] comprised limited sets of images with diverse distortion types, annotated using controlled laboratory studies. Subsequent efforts such as TID2013 [Ponomarenko2015] and MDID [Sun2017] expanded dataset sizes but remained insufficient for training data-hungry deep learning models due to their inherent scale limitations. A key bottleneck is the labor-intensive and costly subjective scoring process, which hinders the creation of large-scale datasets. As a result, only a few sizable datasets such as PDAP-HDDS [Liu2018] exist. Moreover, synthetically marked-up datasets also exist, e.g., KADIS-700k [Lin2020] with markups obtained by a model trained on KADID-10k [Lin2020], but unfortunately their utility is compromised for certain evaluation tasks.
@@
Recent studies [7460955, hosu2018expertise] demonstrate that, with proper experimental design, crowdsourced annotations can achieve a reliability comparable to laboratory settings. Moreover, in prior work [ershov2024reliability], the reliability and stability of using crowdsourcing for stylistic and rendering evaluation was investigated, demonstrating its effectiveness as a viable data collection method. Crowdsourcing is becoming increasingly important and relying on it has enabled the creation of large-scale, diverse IQA datasets such as KonIQ-10k [Hosu2019] and UHD-IQA [Hosu2024], which address the scalability limitations of traditional approaches. Unlike earlier benchmarks constrained by small sample sizes or missing subjective labels, these datasets leverage crowdsourcing to balance scale and annotation quality, making them more suitable for modern data-driven IQA methods.
@@
The aforementioned works primarily address objective IQA, focusing on technical distortions such as noise or compression artifacts. In contrast, a separate line of research targets image aesthetics evaluation, which evaluates subjective preferences and perceptual appeal. Prominent datasets in this domain include AVA [murray2012ava], AVA-PD [kairanbay2018towards], and EVA [kang2020eva]. Unlike technical IQA benchmarks, these datasets evaluate aesthetics by comparing entirely different scenes, where users select the most visually pleasing image within a broad theme, e.g., “flying bird”. Crucially, however, this evaluation captures holistic judgments incorporating content, composition, lighting, and stylistic choices, rather than isolated technical quality. This makes it difficult to assess the contribution of stylization or rendering to overall image quality.
@@
Recently, a growing research trend has emerged to evaluate the quality of AI-generated images, addressing unique challenges such as perceptual realism, semantic consistency, and technical artifacts. Datasets like HPDv2 [wu2023human], AGIQA-3k [Li2023], AIGIQA-20k [li2024aigiqa], and PKU-I2IQA [yuan2023pku] have been developed to assess these aspects. These datasets typically employ a structured methodology: multiple images are generated for each prompt, and then their quality, authenticity, and prompt alignment are rigorously validated using mean opinion scores. This approach provides a standardized framework to assess the rapidly evolving capabilities of generative AI models.
@@
Several tone-mapped image quality datasets have as well been introduced [ak2022rv, jiang2025dataset], in which multiple tone mapping operators (TMOs) are applied to high dynamic range (HDR) images to produce low dynamic range (LDR) outputs of varying perceptual quality (we use the term HDR in the traditional sense of a high-range image destined for reduction to SDR — Standard Dynamic Range — rather than the newer display technology that presents actual HDR to the viewer). While TMOs do not introduce degradations in the traditional sense, making these datasets somewhat aligned with the subjective nature of our work, they differ fundamentally in scope and application. Specifically, the input and output images in such datasets belong to different dynamic ranges, making the transformations more technical than stylistic. Moreover, these datasets do not address or address minimally the broader and more common use cases involving stylistic and rendering variation within the same dynamic range, such as color grading, contrast adjustments, or photographic style emulation, which are the core of EAR and the focus of this work.
@@
The UHD-IQA Benchmark Database [Hosu2024] contains 6,073 ultra-high-definition (4K) images collected from Pixabay and curated to exclude synthetic or heavily retouched content. Image quality was assessed in a single-stimulus protocol, where 10 expert raters composed of photographers and graphic artists judged each image twice on a continuous scale from bad (1%) to excellent (100%), yielding around 20 ratings per image. The resulting mean opinion scores provide ground-truth annotations, with high inter-round consistency measured by the Spearman Rank Order Correlation Coefficient (SRCC), reported at approximately $0.93$. Along with perceptual quality ratings, the dataset includes metadata such as image dimensions, popularity indicators, and machine-generated tags, making it one of the largest and most comprehensive publicly available benchmarks for no-reference image quality assessment in the UHD domain. However, the dataset contains only a single rendering per image, limiting its usefulness for evaluating different renderings of the same scene; moreover, the absence of pairwise comparisons further limits its ability to capture fine-grained perceptual differences.
@@
LIVE [sheikh2005live] 29 779 Distortion 29 $S$
CSIQ [Chandler2010] 30 866 Distortion 25 N/A
VCL@FER [Zari2012] 23 552 Distortion 118 $S$
TID2013 [Ponomarenko2015] 25 3,000 Distortion 971 PC
PDAP-HDDS [Liu2018] 250 12,000 Distortion 38 ACR
KonIQ-10k [Hosu2019] 10,073 0 Distortion+Content 1,459 ACR
KADID-10k [Lin2020] 81 10,125 Distortion 2,209 ACR
UHD-IQA [Hosu2024] 6,073 6,073 Content 10 $S$
@@ 2.2 Benchmark NR-IQA methods
For IQA tasks, numerous methodological approaches exist. The field of IQA is dominated by two primary approaches: full-reference and no-reference methods. Full-reference techniques require a complete, high-fidelity reference image for comparison to quantify perceptual degradation. Conversely, no-reference methods operate without any reference image, directly estimating the quality score based solely on the features of the test image. In our work, since we propose a dataset without reference images (i.e., with no “golden standard” for comparison exists), traditional full-reference metrics like SSIM [wang2004image], PSNR, or LPIPS [zhang2018unreasonable] cannot be directly applied. Consequently, our analysis focuses exclusively on NR-IQA methods.
@@
One of the most widely used statistical approaches for NR-IQA is BRISQUE [Mittal2012]. This method extracts a set of handcrafted features based on natural scene statistics and employs a Support Vector Regressor (SVR) [scholkopf2000new] to predict the final quality score. More recently, PIQUE [pique] has adopted a similar paradigm, leveraging perceptually relevant features in conjunction with a statistical learning framework to estimate image quality.
With the emergence of large-scale IQA datasets, DL-based methods have gained significant traction. Early works such as DeepBIQ [bianco2018use], MUSIQ [ke2021musiq], PaQ-2-PiQ [ying2020patches], CONTRIQUE [madhusudana2022image], Re-IQA [saha2023re], and ARNIQA [arniqa] explore various deep learning paradigms, including patch-based learning, contrastive representation learning, and transformer-based architectures, to improve the accuracy and generalizability of IQA models.
Furthermore, a growing number of recent approaches incorporates vision-language models for image quality assessment. For instance, CLIP-IQA [Wang2023] estimates image quality by computing the cosine similarity between visual embeddings and text prompts such as “This photo is good” and “This photo is bad.” Building upon this idea, LIQE [Zhang2023] improves the framework by fine-tuning both visual and textual encoders and introducing a multi-task optimization strategy, which enhances robustness, particularly due to its use of large-scale training data. Subsequently, QualiCLIP [Agnolucci2024] was designed to be more sensitive to varying levels of image distortion. Most recently, LAR-IQA [lariqa] demonstrated that incorporating KANs [liu2024kan] can further boost model performance.
@@ 3 Proposed DEAR
The MIT-Adobe FiveK dataset [fivek] is a widely-used, high-resolution image dataset designed for research in computational photography and image enhancement. It comprises 5,000 RAW photographs captured by DSLR cameras under diverse lighting and scene conditions. Each image has been independently rendered by five professional photo renderers, labeled Expert A through E, yielding five stylistically distinct reference renders per image. The dataset includes both the original RAW files and their corresponding JPEG renderings, making it suitable for tasks such as supervised learning of image-to-image translation, automatic photo retouching, tone and exposure correction, white balance adjustment, style-specific enhancement, etc. Its diversity in photographic content, expert renderers, and high-resolution format with up to $5616\times 3744$ pixels makes it a valuable resource for training and evaluating deep learning models in both paired and unpaired settings.
The distribution of labels for these images that is available within the MIT-Adobe FiveK dataset is shown in Figure 2, together with another distribution obtained by means of querying the BLIP [li2022blip]. In order to validate the overall quality of BLIP results we randomly selected 100 images from the dataset and marked them up. The average accuracy for all categories equals $0.96$, while F1-score equals $0.86$. More detailed analysis of category markup can be seen in the Supplementary material. The variety of the dataset is sufficient to draw general conclusions.
@@
Each of the MIT-Adobe FiveK images is rendered in six different styles: five photographers, i.e. renderers and one neutral style from Adobe Photoshop auto settings, which yields $\binom{6}{2}=15$ possible unique style pair combinations per image. To perform subjective quality assessment, these pairs were uploaded to the crowdsourcing platform Yandex.Tasks. The Yandex.Tasks interface displayed six image pairs per page. For each pair, the participants, i.e., evaluators were asked: “Which image is preferable?”. Images were presented against a 50% gray background, following widely used recommendations [mantiuk2012comparison] and standard to minimize bias in color perception by using a neutral background. To further ensure unbiased assessment, no additional controls or information were placed between the image pairs. Each image pair was also spaced adequately to simplify the visual comparison. An example of such an image pair can be seen in Figure 3.
@@
The evaluators could select one of the following options for each image pair: “The left image is preferable”, “The right image is preferable” and “Both images are equally preferable”. To identify unreliable evaluators, each page included an additional pair of nearly identical images as a control question. Essentially, the whole voting process was similar to the one in [ershov2024reliability].

### v4

@@ Abstract
Figure 1: Limitations of existing IQA methods: while LIQE [45] assigns nearly identical quality scores with standard deviations from the mean value 3.5% across evaluations, subjective evaluation reveals clear differences. Our work addresses this gap by introducing a new benchmark Dataset for Evaluating of Aesthetics of Rendering (DEAR).
@@ 1 Introduction
Image Quality Assessment refers to the task of automatically evaluating the quality of images in a way that aligns with human perception. Accurate IQA is critical for the development and optimization of computer vision and image processing algorithms, as it provides an objective signal that guides models toward producing results that are visually pleasing or perceptually faithful [43]. Whether in image compression, enhancement, restoration, or generation, IQA metrics are expected to help ensure that algorithmic outputs meet human expectations of quality, as well as to measure the performance of the algorithms themselves.
Traditional IQA research has focused primarily on assessing the degradation of perceived image quality introduced by the presence of distortions such as noise, blur, or compression artifacts. Over time, however, the field recognized that perceived image quality extends beyond simple degradations: viewers judge an image not only by the absence of artifacts, but also by how convincingly it conveys realism and visual appeal. This shift in perspective motivated the creation of modern subjective IQA datasets, such as UHD-IQA [13], which capture human judgments on more holistic attributes, including composition, resolution, and color harmony, without requiring a pristine reference. However, even in subjective IQA datasets [13], participants are typically asked to rate the overall quality of individual images, not to compare different stylistic renditions of the same scene. As a result, such datasets capture judgments about technical fidelity or perceptual realism rather than aesthetics of rendering, and are not suited for modeling user preferences between alternative rendering interpretations, which could be captured much more easily if datasets such as UHD-IQA contained pairwise evaluation results.
Evaluation of aesthetics of rendering introduces a qualitatively different challenge. Aesthetics of rendering is not characterized by degradations, but by aesthetic attributes such as color usage, texture, and artistic expression. Importantly, while image degradations are known to be of lesser quality than the original image, multiple renderings can be equally valid and pleasing, making rendering preference inherently subjective and context-dependent. Unlike quality degradations, which tend to have a broadly agreed-upon negative impact, rendering variation often reflects personal, cultural, or situational preferences. Finally, because there is no objectively correct answer for the most pleasant rendering, it is impossible to guarantee whether users select their responses attentively, making the annotation process vulnerable to careless behavior [11].
@@
To address this gap, we introduce the Dataset for Evaluating the Aesthetics of Rendering (DEAR), a new benchmark specifically designed to allow learning subjective judgments of image rendering. DEAR is constructed by augmenting the MIT-Adobe FiveK dataset [36] with pairwise human preference scores obtained through crowdsourced evaluation platform. Each pair of image renderings was evaluated by at least $25$ evaluators, with a total of $13,648$ unique evaluators participating overall. These annotations enable the development of EAR models that go beyond degradation-focused assessments and instead capture the more nuanced, human-centered perception of rendering.
@@ 2.1 Benchmark IQA datasets
In recent years, numerous IQA datasets have been introduced. Benchmarks such as LIVE [30], VCL@FER [4], NITS-IQA [27], and CSIQ [9] comprised limited sets of images with diverse distortion types, annotated using controlled laboratory studies. Subsequent efforts such as TID2013 [26] and MDID [39] expanded dataset sizes but remained insufficient for training data-hungry deep learning models due to their inherent scale limitations. A key bottleneck is the labor-intensive and costly subjective scoring process, which hinders the creation of large-scale datasets. As a result, only a few sizable datasets such as PDAP-HDDS [32] exist. Moreover, synthetically marked-up datasets also exist, e.g., KADIS-700k [12] with markups obtained by a model trained on KADID-10k [12], but unfortunately their utility is compromised for certain evaluation tasks.
@@
Recent studies [31, 14] demonstrate that, with proper experimental design, crowdsourced annotations can achieve a reliability comparable to laboratory settings. Moreover, in prior work [11], the reliability and stability of using crowdsourcing for stylistic and rendering evaluation was investigated, demonstrating its effectiveness as a viable data collection method. Crowdsourcing is becoming increasingly important and relying on it has enabled the creation of large-scale, diverse IQA datasets such as KonIQ-10k [35] and UHD-IQA [13], which address the scalability limitations of traditional approaches. Unlike earlier benchmarks constrained by small sample sizes or missing subjective labels, these datasets leverage crowdsourcing to balance scale and annotation quality, making them more suitable for modern data-driven IQA methods.
@@
The aforementioned works primarily address objective IQA, focusing on technical distortions such as noise or compression artifacts. In contrast, a separate line of research targets image aesthetics evaluation, which evaluates subjective preferences and perceptual appeal. Prominent datasets in this domain include AVA [24], AVA-PD [16], and EVA [17]. Unlike technical IQA benchmarks, these datasets evaluate aesthetics by comparing entirely different scenes, where users select the most visually pleasing image within a broad theme, e.g., “flying bird”. Crucially, however, this evaluation captures holistic judgments incorporating content, composition, lighting, and stylistic choices, rather than isolated technical quality. This makes it difficult to assess the contribution of stylization or rendering to overall image quality.
@@
Recently, a growing research trend has emerged to evaluate the quality of AI-generated images, addressing unique challenges such as perceptual realism, semantic consistency, and technical artifacts. Datasets like HPDv2 [40], AGIQA-3k [10], AIGIQA-20k [19], and PKU-I2IQA [42] have been developed to assess these aspects. These datasets typically employ a structured methodology: multiple images are generated for each prompt, and then their quality, authenticity, and prompt alignment are rigorously validated using mean opinion scores. This approach provides a standardized framework to assess the rapidly evolving capabilities of generative AI models.
@@
Several tone-mapped image quality datasets have as well been introduced [3, 15], in which multiple tone mapping operators (TMOs) are applied to high dynamic range (HDR) images to produce low dynamic range (LDR) outputs of varying perceptual quality (we use the term HDR in the traditional sense of a high-range image destined for reduction to SDR — Standard Dynamic Range — rather than the newer display technology that presents actual HDR to the viewer). While TMOs do not introduce degradations in the traditional sense, making these datasets somewhat aligned with the subjective nature of our work, they differ fundamentally in scope and application. Specifically, the input and output images in such datasets belong to different dynamic ranges, making the transformations more technical than stylistic. Moreover, these datasets do not address or address minimally the broader and more common use cases involving stylistic and rendering variation within the same dynamic range, such as color grading, contrast adjustments, or photographic style emulation, which are the core of EAR and the focus of this work.
@@
The UHD-IQA Benchmark Database [13] contains 6,073 ultra-high-definition (4K) images collected from Pixabay and curated to exclude synthetic or heavily retouched content. Image quality was assessed in a single-stimulus protocol, where 10 expert raters composed of photographers and graphic artists judged each image twice on a continuous scale from bad (1%) to excellent (100%), yielding around 20 ratings per image. The resulting mean opinion scores provide ground-truth annotations, with high inter-round consistency measured by the Spearman Rank Order Correlation Coefficient (SRCC), reported at approximately $0.93$. Along with perceptual quality ratings, the dataset includes metadata such as image dimensions, popularity indicators, and machine-generated tags, making it one of the largest and most comprehensive publicly available benchmarks for no-reference image quality assessment in the UHD domain. However, the dataset contains only a single rendering per image, limiting its usefulness for evaluating different renderings of the same scene; moreover, the absence of pairwise comparisons further limits its ability to capture fine-grained perceptual differences.
@@
LIVE [30] 29 779 Distortion 29 $S$
CSIQ [9] 30 866 Distortion 25 N/A
VCL@FER [4] 23 552 Distortion 118 $S$
TID2013 [26] 25 3,000 Distortion 971 PC
PDAP-HDDS [32] 250 12,000 Distortion 38 ACR
KonIQ-10k [35] 10,073 0 Distortion+Content 1,459 ACR
KADID-10k [12] 81 10,125 Distortion 2,209 ACR
UHD-IQA [13] 6,073 6,073 Content 10 $S$
@@ 2.2 Benchmark NR-IQA methods
For IQA tasks, numerous methodological approaches exist. The field of IQA is dominated by two primary approaches: full-reference and no-reference methods. Full-reference techniques require a complete, high-fidelity reference image for comparison to quantify perceptual degradation. Conversely, no-reference methods operate without any reference image, directly estimating the quality score based solely on the features of the test image. In our work, since we propose a dataset without reference images (i.e., with no “golden standard” for comparison exists), traditional full-reference metrics like SSIM [38], PSNR, or LPIPS [44] cannot be directly applied. Consequently, our analysis focuses exclusively on NR-IQA methods.
@@
One of the most widely used statistical approaches for NR-IQA is BRISQUE [5]. This method extracts a set of handcrafted features based on natural scene statistics and employs a Support Vector Regressor (SVR) [29] to predict the final quality score. More recently, PIQUE [34] has adopted a similar paradigm, leveraging perceptually relevant features in conjunction with a statistical learning framework to estimate image quality.
With the emergence of large-scale IQA datasets, DL-based methods have gained significant traction. Early works such as DeepBIQ [7], MUSIQ [18], PaQ-2-PiQ [41], CONTRIQUE [22], Re-IQA [28], and ARNIQA [1] explore various deep learning paradigms, including patch-based learning, contrastive representation learning, and transformer-based architectures, to improve the accuracy and generalizability of IQA models.
Furthermore, a growing number of recent approaches incorporates vision-language models for image quality assessment. For instance, CLIP-IQA [37] estimates image quality by computing the cosine similarity between visual embeddings and text prompts such as “This photo is good” and “This photo is bad.” Building upon this idea, LIQE [45] improves the framework by fine-tuning both visual and textual encoders and introducing a multi-task optimization strategy, which enhances robustness, particularly due to its use of large-scale training data. Subsequently, QualiCLIP [2] was designed to be more sensitive to varying levels of image distortion. Most recently, LAR-IQA [6] demonstrated that incorporating KANs [21] can further boost model performance.
@@ 3 Proposed DEAR
The MIT-Adobe FiveK dataset [36] is a widely-used, high-resolution image dataset designed for research in computational photography and image enhancement. It comprises 5,000 RAW photographs captured by DSLR cameras under diverse lighting and scene conditions. Each image has been independently rendered by five professional photo renderers, labeled Expert A through E, yielding five stylistically distinct reference renders per image. The dataset includes both the original RAW files and their corresponding JPEG renderings, making it suitable for tasks such as supervised learning of image-to-image translation, automatic photo retouching, tone and exposure correction, white balance adjustment, style-specific enhancement, etc. Its diversity in photographic content, expert renderers, and high-resolution format with up to $5616\times 3744$ pixels makes it a valuable resource for training and evaluating deep learning models in both paired and unpaired settings.
The distribution of labels for these images that is available within the MIT-Adobe FiveK dataset is shown in Figure 2, together with another distribution obtained by means of querying the BLIP [20]. In order to validate the overall quality of BLIP results we randomly selected 100 images from the dataset and marked them up. The average accuracy for all categories equals $0.96$, while F1-score equals $0.86$. More detailed analysis of category markup can be seen in the Supplementary material. The variety of the dataset is sufficient to draw general conclusions.
@@
Each of the MIT-Adobe FiveK images is rendered in six different styles: five photographers, i.e. renderers and one neutral style from Adobe Photoshop auto settings, which yields $\binom{6}{2}=15$ possible unique style pair combinations per image. To perform subjective quality assessment, these pairs were uploaded to the crowdsourcing platform Yandex.Tasks. The Yandex.Tasks interface displayed six image pairs per page. For each pair, the participants, i.e., evaluators were asked: “Which image is preferable?”. Images were presented against a 50% gray background, following widely used recommendations [23] and standard to minimize bias in color perception by using a neutral background. To further ensure unbiased assessment, no additional controls or information were placed between the image pairs. Each image pair was also spaced adequately to simplify the visual comparison. An example of such an image pair can be seen in Figure 3.
@@
The evaluators could select one of the following options for each image pair: “The left image is preferable”, “The right image is preferable” and “Both images are equally preferable”. To identify unreliable evaluators, each page included an additional pair of nearly identical images as a control question. Essentially, the whole voting process was similar to the one in [11].